OpenStudy (anonymous):

how to write 0.092 as a fraction or mixed number

OpenStudy (anonymous):

the fraction would be: \[\frac{ 92 }{ 1000 }\] The reduced fraction would be: \[\frac{ 23 }{ 25 }\] (because 4 was the dividing factor) and last but not least the mixed no. would be \[1\frac{ 2}{ 25 }\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."